Calculate the probabilities of P(0<z<z0)= 0.1700
1
Expert's answer
2021-09-30T09:34:24-0400

"P(0<z<z_0)=P(z<z_0)-P(z<0)=P(z<z_0)-0.5=0.17"

"P(z<z_0)=0.5+0.17=0.67"

"z_0=0.44"